Check My Work 10-2: Basic Definitions 10-5: The Cost of Retained Earnings, rs Cost of Common Equity and WACc Patton Paints Corporation has a target capital structure of 30% debt and 70% common equity, with no preferred stock. Its before-tax cost of debt is 9% and its marginal tax rate is 40%. The current stock price is Po-$22.00. The last dividend was Do = $2.25, and it is expected to grow at a 8% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. a. rs b. WACC
Answer #1
Calculation of cost of equity
We can calculate the cost of common equity using dividend capitalization model stated as under,
Cost of common equity = [Dividend per share for next year / Current Market value of stock] * Growth rate of dividends
Dividend per share for next year = last dividend * (1+growth rate) = $2.25 * (1+0.08) = 2.43
Cost of common equity = [$2.43 / $22] * 0.08 = 0.190455
Cost of common equity = 19.05%
Calculation of WACC i.e.weighted average cost of capital
WACC = [After tax cost of debt * Weight of debt in capital structure] + [Cost of equity * Weight of equity in capital structure]
After tax cost of debt = Before tax cost of debt * (1 - Tax rate) = 9% * (1-0.40) = 5.40%
WACC = [0.054 * 30%] + [0.1905 * 70%] = 0.1495
WACC = 14.95%